# Yong Liu    Guo Qin    Zhiyuan Shi    Zhi Chen    Caiyin Yang    Xiangdong Huang    Jianmin Wang    Mingsheng Long

### v2

@@ 4.2 TimeBench
We collected and curated TimeBench, which comprises over a trillion time points from various sources, as shown in Figure 3. Several datasets originate from research teams (Woo et al., 2024; Ansari et al., 2024; Liu et al., 2024a, b). While most datasets are collected from real-world records, a small portion (0.05%) is generated synthetically to enhance pattern diversity, following KernelSynth proposed by Ansari et al. (2024). We also leverage substantial meteorological data (Hersbach et al., 2020) because of the predictability of weather systems. Data of different frequencies encompasses common and comprehensive temporal dynamics.
@@ 5 Experiments
Table 1: Zero-shot forecasting results of time series foundation models on long-term forecasting datasets (Time-Series-Library) (Wu et al., 2022). Corresponding prediction lengths include $\{96,192,336,720\}$. A lower MSE or MAE indicates a better prediction. Averaged results of four prediction lengths are reported here. $1^{\text{st}}$ Count represents the number of wins achieved by a model under all prediction lengths and datasets. Results of baseline models are officially reported by Shi et al. (2024b). Datasets in pre-training are not evaluated on corresponding models, which are denoted by the dash (--). Full results under all prediction lengths are provided in Table 9.
@@
[…]
@@
Table 2: GIFT-Eval comprises $23$ datasets characterized by a variety of frequencies, variate numbers, and prediction lengths. We evaluate zero-shot performance using $100$ generated series, being consistent with Woo et al. (2024). A lower MASE or CRPS indicates a better performance. Rank assigns a numerical ranking of all $97$ configurations. Baseline results are officially reported by Aksu et al. (2024).
@@
Naïve ARIMA Theta (2020) (2023a) (2019) (2022) (2023a) (2023b) (2025) (2024) (2024) (Ours)
@@ 5.1 Time Series Forecasting
In this section, we focus on zero-shot forecasting, we compare Sundial with advanced time series foundation models on various benchmarks, including (1) point forecasting: we adopt the long-term forecasting benchmark (Wu et al., 2022), which assesses the performance under different forecasting horizons using MSE and MAE; (2) probabilistic forecasting: we experiment on GIFT-Eval (Aksu et al., 2024) and FEV leaderboard (Ansari et al., 2024), following their official evaluation suite and assessing point (MASE) and probabilistic (CRPS and WQL) metrics. All evaluated datasets are excluded from the pre-training dataset. Model is available on HuggingFace [^1] https://huggingface.co/thuml/sundial-base-128m. and configurations are detailed in Table 5.
@@ 5.1.1 Point Forecasting
As shown in Table 1, Sundial consistently outperforms other advanced time series foundation models. Compared with the previous state-of-the-art model Time-MoE (Shi et al., 2024b), the Sundial family using fewer parameters achieves the average MSE reduction of $7.57\%$ and averaged MAE reduction of $4.71\%$. Notably, continuous tokenization allows our model to conduct patch-level forecasting with fewer autoregression steps, while Chronos using point-wise discrete tokenization may not be suitable in long-term forecasting.
Figure 4: Model evaluation on the FEV leaderboard, which includes $27$ datasets not seen by Sundial. Baseline models can be categorized into statistical methods fitting on each time series, task-specific deep models trained on each dataset, and pre-trained foundation models. Pre-trained Models that have seen several datasets during pre-training are denoted as Pre-trained Models (Other). A lower MASE/WQL indicates a better result. Sundial makes probabilistic predictions using $20$ generated series, being consistent with Ansari et al. (2024).
@@ 5.1.2 Probabilistic Forecasting
Beyond point forecasting, Sundial possesses a unique generative capability for making probabilistic predictions. Following Ansari et al. (2024), we calculate the median and quantiles using a set of raw predictions of Sundial. While several baseline models have been pre-trained by the consistent objective function for probabilistic evaluation, e.g., quantile loss for WQL, Sundial calculates these statistics for evaluation without any prior knowledge.
@@
Aggregated results are presented in Table 2. The benchmark evaluates performance from $23$ datasets and $13$ baseline models, encompassing statistical methods, task-specific models, and time series foundation models. Among supervised models and advanced foundation models, Sundial attains the first place in MASE and second place in CRPS on all unseen datasets. While the top PatchTST (Nie et al., 2022) is exhaustively trained and tweaked on each dataset, the zero-shot performance of Sundial highlights its simplicity and robustness on this comprehensive benchmark.
@@
We evaluate our Sundial on the open leaderboard established by AutoGluon (Ansari et al., 2024), which includes $27$ datasets for probabilistic forecasting. As shown in Figure 4, the zero-shot forecasting performance of Sundial exceeds $70\%$ statistical methods and deep models that are superwisedly trained in distribution. While Sundial is ranked as the second zero-shot pre-trained models after Chronos, Sundial realizes 35×35$\times$35 × inference speedup as shown in Figure 5. Based on patch-wise tokenization and multi-patch prediction, our inference speed is near to N-BEATS.
@@
Figure 5: Inference time evaluation following Ansari et al. (2024), which is averaged from the FEV leaderboard. Computing resources of different models are marked. We plot the logarithmic x-axis.
@@ 5.3 TimeFlow Loss
Based on the flow-matching framework, TimeFlow Loss allows autoregressive models to learn and generate flexible distributions while enhancing representation learning. To validate the effectiveness of this design, we implement two alternatives: (1) an MLP network and MSE Loss and (2) a parameterized training objective based on the denoising diffusion procedure (Li et al., 2024). We adopt the same parameterized network and Transformer backbone and pre-train them on TimeBench. Since the converged training loss is not comparable across different objective functions, we compare zero-shot performance in Table 3. Despite allowing for sampling predictions, performance using diffusion-based objective is notably inferior to TimeFlow Loss.
@@ 5.5 Model Adaptation
Inspired by the prevalence of instruction tuning (Wei et al., 2021) that adapts foundation models on a collection of tasks. We fine-tune pre-trained Sundial (Base) on the FEV leaderboard, including short-term tasks with different prediction lengths. Our model is tuned once on all aggregated datasets. We evaluate the performance on unseen test splits (Figure 8). We observe that the performance can be further improved compared to zero-shot forecasting. Furthermore, training from scratch on aggregated datasets results in inferior performance, implying knowledge transfer in pre-trained models.
@@
Figure 9: Ablation studies with respect to architectural enhancements. We report the averaged results of TSLib datasets (Wu et al., 2022) from four prediction lengths $\{96,192,336,720\}$ and all six datasets. The context length is set to $2880$ and the patch length is $16$.
@@ 5.6 Ablation Study
Prior research (Liu et al., 2024a) observed that the introduction of RoPE (Su et al., 2024) yields better results in supervised forecasting tasks. As shown in Figure 9 (a), RoPE can also improve zero-shot forecasting, presenting a general enhancement for time series foundation models.
@@
Pre-LN (Baevski & Auli, 2018) is widely adopted in large language models (Touvron et al., 2023) due to the training stability. As depicted in Figure 9 (b), training with Pre-LN for more iterations yield better performance. In contrast, training with Post-LN, which is the predominant choice in supervised models, may adversely affect downstream results.
@@
We make it to leverage FlashAttention (Dao et al., 2022) and KV Cache to reduce the computational costs. As shown in Figure 9 (c) and (d), they notably reduce $14.8\%$ memory footprint and $43.6\%$ inference time without affecting performance.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (U2342217 and 62021002), the BNRist Innovation Fund (BNR2024RC01010), and the National Engineering Research Center for Big Data Software.
@@
This paper aims to advance the development of time series foundation models. We curated a pre-training dataset from publicly available resources. Our models employ an efficient tokenization and incorporate a generative training objective. The proposed TimeFlow Loss provides insights for training generative foundation models for time series forecasting. We released our pre-trained models that demonstrate notable zero-shot forecasting performance and versatility across different benchmarks. The generative forecasting paradigm enhances the model reliability for decision-making. The millisecond inference speed of our model can be beneficial to many practitioners. Our paper mainly focuses on scientific research and has no obvious negative social impact.
@@ Appendix A Dataset Statistics
Large-scale datasets are of paramount importance for pre-training foundation models. Recent research has contributed significant time series datasets (Das et al., 2023b; Liu et al., 2024b; Shi et al., 2024b). While the scaling law of time series foundation models has been explored in the recent work (Shi et al., 2024a), the pre-training scale remains relatively limited. Given the heterogeneity of time series compared to other modalities, it raises the question of whether it is feasible to learn from enormous series. To address the question, we curated TimeBench with a trillion time points from various domains.
Unlike other modalities, most time series are unavailable on open websites or repositories. There are also limited domains that encompass typical and predictable time series, leading to slow progress on dataset construction. Therefore, we conducted tedious preprocessing, including missing values imputation, abnormalities exclusion, and normalization techniques. We conducted statistical analysis, examining time series through the lenses of intrinsic properties, e.g., non-stationarity, forecastability, and seasonality. This approach allows us to characterize the data quality inherent to time series, which affects the training stability of next-token prediction. We also adopt synthetic techniques to improve pattern diversity. Further, we adopt ERA5 (Muñoz-Sabater et al., 2021), including systematic real-world temporal observations.
The statistical details of TimeBench are summarized in Table 4. In addition to open-source datasets from research teams on time series foundation models (Woo et al., 2024; Ansari et al., 2024; Liu et al., 2024b, a), we collected substantial real-world time series from various domains such as finance, IoT, meteorology, and healthcare (Goldberger et al., 2000). These resources enable us to construct large-scale time-series corpora exceeding a trillion time points. The corpora include highly credible and predictable data with a wide range of frequencies, lengths, and numbers of variates, providing comprehensive temporal dynamics and variation patterns to facilitate downstream applications. To prevent data leakage, we exclude all datasets evaluated in Section 5.1 to make sure that Sundial conducts zero-shot forecasting.
@@
(2024) (2000) (Ours) (Ours) (2024) (2024) (2021) (2021) (2021) (2021) (2021) (2021)
@@ Appendix B Implementation Details
All experiments are implemented using PyTorch (Paszke et al., 2019) and executed with $32$ NVIDIA A100 GPUs. We employ the AdamW optimizer (Kingma & Ba, 2014) for model optimization. We adopt S3 format (Liu et al., 2024b) for univariate pre-training. During training, data from different domains is sampled according to a predefined ratio to balance the domain weightings and ensure diversity in the training data. We implement a global shuffle strategy by loading time series into a standard parquet format. We use variable-wise normalization to unify the scope of values.
On the FEV leaderboard (Ansari et al., 2024), which consists of short-term forecasting datasets, we train Sundial models by TimeFlow Loss with the prediction length of $F=16$. For the point forecasting (Wu et al., 2022) and GIFT-Eval (Aksu et al., 2024), which consist of forecasting datasets with a prediction length ranging from $6$ to $900$, we train Sundial models by TimeFlow Loss with the prediction length of $F=720$. For the required prediction length less than the model prediction length, we truncate the output generated by Sundial. For the required length more than the prediction horizon, we conduct rolling forecasting. Following Chronos (Ansari et al., 2024), we sample $20$ raw predictions to calculate MASE and WQL on FEV. Being consistent to Moirai (Woo et al., 2024), we sample $100$ raw predictions to calculate MASE and CRPS for GIFT-Eval. The sampling step is fixed as $K=50$. Configurations of Sundial in different sizes are provided in Table 5. We provide a model summary in Table 6, which summarizes several aspects of current time series foundation models.
@@
[…]
$\textbf{Sundial}_{\textit{Small}}$ $16$ $2880$$\{16,720\}$ 6 $(512,2048)$ $8$ $(512,3)$ $32$M
$\textbf{Sundial}_{\textit{Base}}$ $16$ $2880$$\{16,720\}$ 12 $(768,3072)$ $12$ $(768,3)$ $128$M
$\textbf{Sundial}_{\textit{Large}}$ $16$ $2880$$\{16,720\}$ 24 $(1024,4096)$ $16$ $(1024,6)$ $444$M
@@
(Ours) (2024b) (2024a) (2024) (2024) (2024) (2024) (2023) (2023b)
@@ C.1 Discussion of Mode Collapse
Our work addresses this phenomenon through generative modeling. Generative forecasters learn flexible distributions without relying on probabilistic priors. We evaluate the distributional metric Continuous Ranked Probability Score (CRPS) to assess the quality of generated predictions across different training objectives. The results indicate that the predictive distribution modeled by TimeFlow is more coherent and diverse compared to alternative training objectives, particularly on the highly diverse GIFT-Eval (Aksu et al., 2024). It validates the effectiveness of TimeFlow in mitigating mode collapse.

### v3

@@ 4.2 TimeBench
We collected and curated TimeBench, which comprises over a trillion time points from various sources, as shown in Figure 3. Several datasets originate from research teams (woo2024unified; ansari2024chronos; liu2024timer; liutimer). While most datasets are collected from real-world records, a small portion (0.05%) is generated synthetically to enhance pattern diversity, following KernelSynth proposed by ansari2024chronos. We also leverage substantial meteorological data (hersbach2020era5) because of the predictability of weather systems. Data of different frequencies encompasses common and comprehensive temporal dynamics.
@@ 5 Experiments
Table 1: Zero-shot forecasting results of time series foundation models on long-term forecasting datasets (Time-Series-Library) (wu2022timesnet). Corresponding prediction lengths include $\{96,192,336,720\}$. A lower MSE or MAE indicates a better prediction. Averaged results of four prediction lengths are reported here. $1^{\text{st}}$ Count represents the number of wins achieved by a model under all prediction lengths and datasets. Results of baseline models are officially reported by shi2024time. Datasets in pre-training are not evaluated on corresponding models, which are denoted by the dash (-). Full results under all prediction lengths are provided in Table 9.
@@
$\textbf{Time-MoE}_{\textit{Base}}$
$\textbf{Time-MoE}_{\textit{Large}}$
$\textbf{Time-MoE}_{\textit{Ultra}}$
@@
Table 2: GIFT-Eval comprises $23$ datasets characterized by a variety of frequencies, variate numbers, and prediction lengths. We evaluate zero-shot performance using $100$ generated series, being consistent with woo2024unified. A lower MASE or CRPS indicates a better performance. Rank assigns a numerical ranking of all $97$ configurations. Baseline results are officially reported by aksugift.
@@
Naïve ARIMA Theta (salinas2020deepar) (das2023long) (oreshkin2019n) (nie2022time) (liu2023itransformer) (das2023decoder) (hoo2025tabular) (ansari2024chronos) (woo2024unified) (Ours)
@@ 5.1 Time Series Forecasting
In this section, we focus on zero-shot forecasting, we compare Sundial with advanced time series foundation models on various benchmarks, including (1) point forecasting: we adopt the long-term forecasting benchmark (wu2022timesnet), which assesses the performance under different forecasting horizons using MSE and MAE; (2) probabilistic forecasting: we experiment on GIFT-Eval (aksugift) and FEV leaderboard (ansari2024chronos), following their official evaluation suite and assessing point (MASE) and probabilistic (CRPS and WQL) metrics. All evaluated datasets are excluded from the pre-training dataset. Model is available on HuggingFace [^1] https://huggingface.co/thuml/sundial-base-128m. and configurations are detailed in Table 5.
@@ 5.1.1 Point Forecasting
As shown in Table 1, Sundial consistently outperforms other advanced time series foundation models. Compared with the previous state-of-the-art model Time-MoE (shi2024time), the Sundial family using fewer parameters achieves the average MSE reduction of $7.57\%$ and averaged MAE reduction of $4.71\%$. Notably, continuous tokenization allows our model to conduct patch-level forecasting with fewer autoregression steps, while Chronos using point-wise discrete tokenization may not be suitable in long-term forecasting.
Figure 4: Model evaluation on the FEV leaderboard, which includes $27$ datasets not seen by Sundial. Baseline models can be categorized into statistical methods fitting on each time series, task-specific deep models trained on each dataset, and pre-trained foundation models. Pre-trained Models that have seen several datasets during pre-training are denoted as Pre-trained Models (Other). A lower MASE/WQL indicates a better result. Sundial makes probabilistic predictions using 2020 generated series, being consistent with ansari2024chronos.
@@ 5.1.2 Probabilistic Forecasting
Beyond point forecasting, Sundial possesses a unique generative capability for making probabilistic predictions. Following ansari2024chronos, we calculate the median and quantiles using a set of raw predictions of Sundial. While several baseline models have been pre-trained by the consistent objective function for probabilistic evaluation, e.g., quantile loss for WQL, Sundial calculates these statistics for evaluation without any prior knowledge.
@@
Aggregated results are presented in Table 2. The benchmark evaluates performance from $23$ datasets and $13$ baseline models, encompassing statistical methods, task-specific models, and time series foundation models. Among supervised models and advanced foundation models, Sundial attains the first place in MASE and second place in CRPS on all unseen datasets. While the top PatchTST (nie2022time) is exhaustively trained and tweaked on each dataset, the zero-shot performance of Sundial highlights its simplicity and robustness on this comprehensive benchmark.
@@
We evaluate our Sundial on the open leaderboard established by AutoGluon (ansari2024chronos), which includes $27$ datasets for probabilistic forecasting. As shown in Figure 4, the zero-shot forecasting performance of Sundial exceeds $70\%$ statistical methods and deep models that are superwisedly trained in distribution. While Sundial is ranked as the second zero-shot pre-trained models after Chronos, Sundial realizes $35\times$ inference speedup as shown in Figure 5. Based on patch-wise tokenization and multi-patch prediction, our inference speed is near to N-BEATS.
@@
Figure 5: Inference time evaluation following ansari2024chronos, which is averaged from the FEV leaderboard. Computing resources of different models are marked. We plot the logarithmic x-axis.
@@ 5.3 TimeFlow Loss
Based on the flow-matching framework, TimeFlow Loss allows autoregressive models to learn and generate flexible distributions while enhancing representation learning. To validate the effectiveness of this design, we implement two alternatives: (1) an MLP network and MSE Loss and (2) a parameterized training objective based on the denoising diffusion procedure (li2024autoregressive). We adopt the same parameterized network and Transformer backbone and pre-train them on TimeBench. Since the converged training loss is not comparable across different objective functions, we compare zero-shot performance in Table 3. Despite allowing for sampling predictions, performance using diffusion-based objective is notably inferior to TimeFlow Loss.
@@ 5.5 Model Adaptation
Inspired by the prevalence of instruction tuning (wei2021finetuned) that adapts foundation models on a collection of tasks. We fine-tune pre-trained Sundial (Base) on the FEV leaderboard, including short-term tasks with different prediction lengths. Our model is tuned once on all aggregated datasets. We evaluate the performance on unseen test splits (Figure 8). We observe that the performance can be further improved compared to zero-shot forecasting. Furthermore, training from scratch on aggregated datasets results in inferior performance, implying knowledge transfer in pre-trained models.
@@
Figure 9: Ablation studies with respect to architectural enhancements. We report the averaged results of TSLib datasets (wu2022timesnet) from four prediction lengths $\{96,192,336,720\}$ and all six datasets. The context length is set to $2880$ and the patch length is $16$.
@@ 5.6 Ablation Study
Prior research (liu2024timer) observed that the introduction of RoPE (su2024roformer) yields better results in supervised forecasting tasks. As shown in Figure 9 (a), RoPE can also improve zero-shot forecasting, presenting a general enhancement for time series foundation models.
@@
Pre-LN (baevski2018adaptive) is widely adopted in large language models (touvron2023llama) due to the training stability. As depicted in Figure 9 (b), training with Pre-LN for more iterations yield better performance. In contrast, training with Post-LN, which is the predominant choice in supervised models, may adversely affect downstream results.
@@
We make it to leverage FlashAttention (dao2022flashattention) and KV Cache to reduce the computational costs. As shown in Figure 9 (c) and (d), they notably reduce $14.8\%$ memory footprint and $43.6\%$ inference time without affecting performance.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (U2342217 and 62021002), State Grid Ningxia Electric Power Co. Science and Technology Project (SGNXYX00SCJS2400058), and the BNRist Innovation Fund (BNR2024RC01010), the National Engineering Research Center for Big Data Software.
@@
This paper aims to advance the development of time series foundation models. We curated a pre-training dataset from publicly available resources. Our models employ an efficient tokenization and incorporate a generative training objective. The proposed TimeFlow Loss provides insights for training generative foundation models for time series forecasting. We released our pre-trained models that demonstrate notable zero-shot forecasting performance. The generative forecasting paradigm enhances the model reliability for decision-making. Our paper mainly focuses on scientific research and has no obvious negative social impact.
@@ Appendix A Dataset Statistics
Large-scale datasets are of paramount importance for pre-training foundation models. Recent research has contributed significant time series datasets (das2023decoder; liutimer; shi2024time). While the scaling law of time series foundation models has been explored in the recent work (shi2024scaling), the pre-training scale remains relatively limited. Given the heterogeneity of time series compared to other modalities, it raises the question of whether it is feasible to learn from enormous series. To address the question, we curated TimeBench with a trillion time points from various domains.
Unlike other modalities, most time series are unavailable on open websites or repositories. There are also limited domains that encompass typical and predictable time series, leading to slow progress on dataset construction. Therefore, we conducted tedious preprocessing, including missing values imputation, abnormalities exclusion, and normalization techniques. We conducted statistical analysis, examining time series through the lenses of intrinsic properties, e.g., non-stationarity, forecastability, and seasonality. This approach allows us to characterize the data quality inherent to time series, which affects the training stability of next-token prediction. We also adopt synthetic techniques to improve pattern diversity. Further, we adopt ERA5 (munoz2021era5), including systematic real-world temporal observations.
The statistical details of TimeBench are summarized in Table 4. In addition to open-source datasets from research teams on time series foundation models (woo2024unified; ansari2024chronos; liutimer; liu2024timer), we collected substantial real-world time series from various domains such as finance, IoT, meteorology, and healthcare (goldberger2000physiobank). These resources enable us to construct large-scale time-series corpora exceeding a trillion time points. The corpora include highly credible and predictable data with a wide range of frequencies, lengths, and numbers of variates, providing comprehensive temporal dynamics and variation patterns to facilitate downstream applications. To prevent data leakage, we exclude all datasets evaluated in Section 5.1 to make sure that Sundial conducts zero-shot forecasting.
@@
(ansari2024chronos) (goldberger2000physiobank) (Ours) (Ours) (woo2024unified) (ansari2024chronos) (munoz2021era5) (munoz2021era5) (munoz2021era5) (munoz2021era5) (munoz2021era5) (munoz2021era5)
@@ Appendix B Implementation Details
All experiments are implemented using PyTorch (paszke2019pytorch) and executed with $32$ NVIDIA A100 GPUs. We employ the AdamW optimizer (kingma2014adam) for model optimization. We adopt S3 format (liutimer) for univariate pre-training. During training, data from different domains is sampled according to a predefined ratio to balance the domain weightings and ensure diversity in the training data. We implement a global shuffle strategy by loading time series into a standard parquet format. We use variable-wise normalization to unify the scope of values.
On the FEV leaderboard (ansari2024chronos), which consists of short-term forecasting datasets, we train Sundial models by TimeFlow Loss with the prediction length of $F=16$. For the point forecasting (wu2022timesnet) and GIFT-Eval (aksugift), which consist of forecasting datasets with a prediction length ranging from $6$ to $900$, we train Sundial models by TimeFlow Loss with the prediction length of $F=720$. For the required prediction length less than the model prediction length, we truncate the output generated by Sundial. For the required length more than the prediction horizon, we conduct rolling forecasting. Following Chronos (ansari2024chronos), we sample $20$ raw predictions to calculate MASE and WQL on FEV. Being consistent to Moirai (woo2024unified), we sample $100$ raw predictions to calculate MASE and CRPS for GIFT-Eval. The sampling step is fixed as $K=50$. Configurations of Sundial in different sizes are provided in Table 5. We provide a model summary in Table 6, which summarizes several aspects of current time series foundation models.
@@
$(P)$ $(T)$ $(F)$ $(L)$ $(D,D_{\text{ff}})$ $H$ $(D_{\text{tf}},L_{\text{tf}})$ $\text{Count}$
$\textbf{Sundial}_{\textit{Small}}$ $16$ $2880$ $\{16,720\}$ 6 $(512,2048)$ 88 $(512,3)$ 3232M
$\textbf{Sundial}_{\textit{Base}}$ $16$ $2880$ $\{16,720\}$ 12 $(768,3072)$ $12$ $(768,3)$ 128128M
$\textbf{Sundial}_{\textit{Large}}$ $16$ $2880$ $\{16,720\}$ 24 $(1024,4096)$ $16$ $(1024,6)$  [^44] M
@@
(Ours) (shi2024time) (liu2024timer) (woo2024unified) (goswami2024moment) (gruver2024large) (ansari2024chronos) (rasul2023lag) (das2023decoder)
@@ C.1 Discussion of Mode Collapse
Our work addresses this phenomenon through generative modeling. Generative forecasters learn flexible distributions without relying on probabilistic priors. We evaluate the distributional metric Continuous Ranked Probability Score (CRPS) to assess the quality of generated predictions across different training objectives. The results indicate that the predictive distribution modeled by TimeFlow is more coherent and diverse compared to alternative training objectives, particularly on the highly diverse GIFT-Eval (aksugift). It validates the effectiveness of TimeFlow in mitigating mode collapse.